# Monodromy in the space of symmetric cubic surfaces with a line

### v1

@@ 1. Introduction
The Galois group of an enumerative problem is a classical object of study in enumerative algebraic geometry. It was first introduced by Jordan as one of the main subjects of interest at the genesis of Galois theory [Jor70]. This idea enjoyed a revival a century later when Harris proved that the Galois group of an enumerative problem agrees with the monodromy group of its associated cover [Har79]. In modern mathematics Galois groups can be approached from a wide number of perspectives, from Hodge theory and hyperbolic geometry [ACT02], to Lie theory [Man06], to numerical analysis and homotopy continuation [LS09], to name a few. [^1] For a lovely introduction to the history and appearance of Galois groups in enumerative geometry, we refer the reader to [SY21].
@@ 1.1. Main results
The (disconnected) symmetric $27$ lines cover $\widetilde{\mathcal{S}}\to\mathcal{S}$ has monodromy group isomorphic to the Klein $4$-group $K_{4}<W(E_{6})$.
The action of $K_{4}$ on $27$ labeled lines is explicitly worked out as permutations in data A.4. This allows us to completely characterize the covering space $\widetilde{\mathcal{S}}$ — it has 12 connected components, with each one corresponding to an explicit $K_{4}$-set; see 5.16 for details.
There are a few reasons why Theorem 1.2 is surprising. First off, the symmetric group $S_{4}$ and symmetric monodromy group $K_{4}$, thought of as subgroups of $W(E_{6})$, intersect trivially — this means that if we want to witness the $S_{4}$-action on a symmetric cubic surface through monodromy, we must leave the symmetric locus in the total moduli space. Second, a number of restrictions coming from Hodge theory constrain the monodromy group to a group of order 96. However, these restrictions provably do not suffice, as we can name explicit elements of this restricted subgroup that cannot arise via symmetric monodromy. This stands in direct contrast with reasoning used when studying similar problems, such as in [ACT10, Section 8].
When the philosophy of “big monodromy” fails to pin down the Galois group given only Hodge-theoretic and geometric constraints, we are tempted to conclude that a counterexample to big monodromy has been found. However in conversations with A. Landesman and N. Salter, we have been made aware of the catch-22 of big monodromy — that it never fails. This leads naturally to the following question:
Question 1.3.
Can further constraints on the Galois group of lines on symmetric cubic surfaces be seen through the lens of Hodge theory? More broadly, are there any restrictions on (equivariant) monodromy problems that are invisible to Hodge theory?
We do not know an answer to either question, only that additional tools were needed in order to make the required restrictions in this symmetric enumerative problem.
@@ 1.2. Paper structure
In Section 2, we review the construction of the (marked) moduli space of cubic surfaces, and explicitly realize the moduli space of symmetric cubic surfaces as a GIT quotient. In Section 3 we give an overview of the work of Allcock–Carlson–Toledo, including fundamental facts about framed cubic surfaces and their associated period maps, which yields their main theorem, a uniformization of the moduli of cubic surfaces by complex hyperbolic $4$-space. In Section 4, we analogously define a period map for symmetric cubic surfaces via symmetric framings of cubic surfaces, and uniformize the moduli space of symmetric cubic surfaces by the complex hyperbolic line. This section finishes with a Hodge theoretic restriction of symmetric monodromy to a group of order $96$, which it turns out will properly contain the symmetric monodromy group. In Section 5 we determine that the symmetric monodromy group is the Klein $4$-group $K_{4}$, and we describe how it acts on the $27$ lines. Moreover, we show that the $27$ lines cover over the symmetric locus $\widetilde{S}\to\mathcal{S}$ splits into $12$ connected components, and explicitly determine the $K_{4}$-set structure associated to each connected component. Finally, Section 6 describes how to alternatively witness the symmetric monodromy group $K_{4}$ in $W(E_{6})$ and $\operatorname{PO}(4,1,\mathbb{F}_{3})$ via representation theoretic constructions. Appendix A contains explicit data regarding how the symmetric group $S_{4}$ and the generators of the symmetric monodromy group $K_{4}$ act on the $27$ lines on the Fermat cubic surface.
@@ 1.3. Acknowledgements
We thank Benson Farb, Frank Sottile, and Jesse Wolfson for their interest and all independently asking us what the monodromy group is for the problem of finding lines on symmetric cubic surfaces. We thank Taylor Brysiewicz and Kisun Lee for their help with homotopy continuation and certified tracking software. Thank you to Aaron Landesman and Nick Salter for sharing their insights regarding the relationship between this work and various aspects of big monodromy problems. Lastly, we thank Daniel Allcock for his interest and elucidating answers regarding various group theoretic aspects of his work with James A. Carlson and Domingo Toledo.
@@ 2.1. Parameter space of cubic surfaces
There is a linear action of $\operatorname{SL}(4,\mathbb{C})$ on $\mathcal{W}^{\mathrm{sm}}$ given by $g\cdot f:=f\circ g^{-1}$. This induces a left action of $\operatorname{PGL}(4,\mathbb{C})$ on the projectivization $\mathbb{P}(\mathcal{W})$.
@@
The action of $\operatorname{SL}(4,\mathbb{C})$ on each of these loci extends to an action of $\operatorname{PGL}(4,\mathbb{C})$ on their projectivizations. We can take the respective GIT quotients to construct various moduli spaces of cubic surfaces.
@@
When we write a moduli space without a subscript, e.g. $\mathcal{M}$, we implicitly mean the moduli of smooth objects.
@@ 2.2. Marked parameter space of cubic surfaces
It is also classically known that the moduli space of marked cubic surfaces $\widehat{\mathcal{M}}$ is isomorphic to the moduli space of cubic surfaces equipped with a line $\widetilde{\mathcal{M}}$ [Bea09, pg. 19]. We shall freely identify these spaces.
@@ 2.3. Parameter space of symmetric cubic surfaces
We see that the the parameter space of symmetric cubic forms $\mathbb{P}(\mathcal{V})=\mathbb{P}^{2}$ embeds linearly into the parameter space of cubic forms $\mathbb{P}(\mathcal{W})$. Define $\Delta^{S_{4}}$ to be the symmetric discriminant curve, which corresponds to the locus of $S_{4}$-invariant singular cubic forms in the parameter space $\mathbb{P}\mathcal{V}$.
@@
The GIT quotient $\mathcal{S}=N_{\operatorname{PGL}(4,\mathbb{C})}(S_{4})\backslash\!\!\backslash\mathbb{P}\mathcal{V}^{\mathrm{sm}}$ is the moduli space of smooth symmetric cubic surfaces.
@@
Suppose that two symmetric cubic forms $f_{1},f_{2}\in\mathbb{P}\mathcal{V}^{\mathrm{sm}}$ determine isomorphic symmetric cubic surfaces $X=Z(f_{1})$ and $Y=Z(f_{2})$. Since such an isomorphism $\varphi:X\xrightarrow[]{\sim}Y$ of varieties preserves their respective canonical classes $K_{X}$ and $K_{Y}$, the map extends to respect their anticanonical embeddings into $\mathbb{P}^{3}$. Thus such an isomorphism $\varphi$ must be the restriction of a linear automorphism coming from the ambient projective space $\mathbb{P}^{3}$. Moreover, the automorphism groups of $X$ and $Y$ must be preserved under such an isomorphism, and so the $S_{4}$-action on $X$ must be sent to the $S_{4}$-action on $Y$. Thus two symmetric cubic surfaces are projectively equivalent when their symmetric cubic forms differ by an element of the normalizer of $S_{4}<\operatorname{PGL}(4,\mathbb{C})$, which was explicitly calculated in 2.10. ∎
@@ 2.4. Symmetry and stability
The inclusion $\mathcal{V}\to\mathcal{W}$ of parameter spaces descends to a inclusion $\iota:\mathcal{S}\to\mathcal{M}$ of the moduli of symmetric cubic surfaces into the whole moduli space. The pullback $\iota^{*}\widehat{\mathcal{M}}$ is then the moduli space of smooth marked symmetric cubic surfaces. Then just as before, the pullback construction let’s us conclude that $\iota^{*}\widehat{\mathcal{M}}$ is isomorphic to the moduli space of symmetric cubic surfaces equipped with a line $\widetilde{\mathcal{S}}$.
@@ 3.4. The period mapping
By 3.3, we have that $\hat{\Gamma}/\hat{\Gamma}^{\prime}\cong W(E_{6})$, and so the cover of $\mathcal{M}$ that $\hat{\Gamma}^{\prime}<\hat{\Gamma}$ corresponds to is the space of cubic surfaces equipped with a line $\widetilde{\mathcal{M}}$. Once we have analogously uniformized moduli space the symmetric cubic surfaces, a special symmetric subgroup of $\mathrm{PO}(4,1,\mathbb{F}_{3})$ will be determined that will contain the monodromy group of the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$.
@@ 4.3. The global period map
There is an isomorphism of groups $N_{\mathrm{U}(4,1)}(S_{4})\cong\mathrm{U}(1,1)\times(\mathrm{U}(1)\times S_{4})$.
@@
Schur’s lemma tells us that the $S_{4}$-centralizer acts by scalars on $W$, and thus is isomorphic to a copy of $\mathrm{U}(1)$ acting on $W$. On the signature $(1,1)$-factor, the $S_{4}$-action is trivial, and thus every element of $\mathrm{U}(1,1)$ arises at an automorphism of the representation $\mathbb{C}\oplus\mathbb{C}$. Thus the normalizer is the product of the normalizers on each factor, proving that $N_{\mathrm{U}(4,1)}(S_{4})\cong\mathrm{U}(1,1)\times(\mathrm{U}(1)\times S_{4})$. ∎
@@
There is an isomorphism of groups $N_{\mathrm{U}(4)}(S_{4})\cong\mathrm{U}(1)\times S_{4}$, where $\mathrm{U}(1)$ acts on the permutation representation $V$ by scalars.
@@
The group $\Gamma=N_{\mathrm{U}(4,1,\mathbb{Z}[\omega])}(S_{4})$ is naturally an arithmetic subgroup of $N_{\mathrm{U}(4,1)}(S_{4})$. Moreover, we have an isomorphism $\Gamma\cong\operatorname{Aut}(\mathrm{diag}(4,-1),\mathbb{Z}[\omega])\times(\langle-\omega\rangle\times S_{4})$
@@
Let $\hat{G}=\mathrm{U}(4,1)$, $\hat{K}=\mathrm{U}(4)$, and $\hat{\Gamma}=\mathrm{U}(4,1,\mathbb{Z}[\omega])$. Set $G=N_{\hat{G}}(S_{4})$, $K=N_{\hat{K}}(S_{4})$, and $\Gamma=N_{\hat{\Gamma}}(S_{4})$. The holomorphic embedding $G/K\to\hat{G}/\hat{K}$ descends to a generically injective finite normalization
@@
Now that we have successfully uniformized the moduli space of symmetric cubic surfaces, we will begin our study of the monodromy group associated to the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$, where $\widetilde{S}$ is the moduli of symmetric cubic surfaces equipped with a line.
@@
One may be tempted to conclude that the monodromy group of the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$ is $K_{4}\times S_{4}$. Indeed, [ACT10, Section 8] outlines why the monodromy groups associated to connected components of moduli of real cubic surfaces are the image of their fundamental groups in $\operatorname{PO}(4,1,\mathbb{F}_{3})$. However, all that 4.12 guarantees is that the monodromy group is contained in $K_{4}\times S_{4}$. Remarkably, the philosophy of “big monodromy” fails to pin down our desired Galois group from purely Hodge theoretic considerations — it will be a proper subgroup of $K_{4}\times S_{4}$! Further analysis using equivariant line geometry on cubic surfaces is required, which we carry out in the next section.

### v2

@@ 1. Introduction
The Galois group of an enumerative problem is a classical object of study in enumerative algebraic geometry. It was first introduced by Jordan as one of the main subjects of interest at the genesis of Galois theory [Jor70]. This idea enjoyed a revival a century later when Harris proved that the Galois group of an enumerative problem agrees with the monodromy group of its associated cover [Har79]. In modern mathematics, Galois groups can be approached from a wide number of perspectives, from Hodge theory and hyperbolic geometry [ACT02], to Lie theory [Man06], to numerical analysis and homotopy continuation [LS09], to name a few. [^1] For a lovely introduction to the history and appearance of Galois groups in enumerative geometry, we refer the reader to [SY21].
@@ 1.1. Main results
The (disconnected) symmetric $27$ lines cover $\widetilde{\mathcal{S}}\to\mathcal{S}$ of moduli spaces has monodromy group isomorphic to $S_{4}\times K_{4}<W(E_{6})$.
In the classical algebraic geometry, deformations and moduli problems were often studied using families of varieties with a chosen projective embedding; we refer to these universal families as parameter spaces, from which the moduli spaces we are interested in can be obtained as a quotient by a projective linear action. For cubic surfaces, the monodromy groups of the 27 lines cover are identical for both the parameter and moduli spaces. In the symmetric locus, we prove that these two monodromy problems diverge.
Let $\mathcal{X}$ denote the moduli space of anti-canonically embedded smooth cubic surfaces and $\mathcal{Y}$ the moduli space of anti-canonically embedded symmetric cubic surfaces (see Section 2 for more detail on how to build these spaces). Stack quotients, in the orbifold sense, of these spaces (by $\operatorname{PGL}_{4}$ and a group $N$ respectively, see Section 2 for details) yield the moduli spaces $\mathcal{M}$ and $\mathcal{S}$. Each space admits a 27 lines cover ($\widetilde{\mathcal{X}}$ and $\widetilde{\mathcal{Y}}$, respectively) which each fits into a commutative diagram with deck groups given like so:
[…]
We determine the monodromy group $M$ which appears in the diagram above:
Theorem 1.3.
The (disconnected) symmetric $27$ lines cover of parameter spaces $\widetilde{\mathcal{Y}}\to\mathcal{Y}$ has monodromy group $M$ isomorphic to centralizer $C_{W(E_{6})}(S_{4})$, which is a Klein 4-group.
The action of $K_{4}$ on $27$ labeled lines is explicitly worked out as permutations in data A.4. This allows us to completely characterize the covering space $\widetilde{\mathcal{Y}}$ — it has 12 connected components, with each one corresponding to an explicit $K_{4}$-set; see 5.13 for details.
There are a few reasons why Theorem 1.3 is interesting. First off, the symmetric group $S_{4}$ and symmetric monodromy group $K_{4}$, thought of as subgroups of $W(E_{6})$, intersect trivially — this means that if we want to witness the $S_{4}$-action on a symmetric cubic surface through monodromy, we must leave the symmetric locus in the total parameter space. Second, the restrictions coming from Hodge theory constrain the monodromy group to a group of order 96 (this is Theorem 1.2). However, these restrictions provably do not suffice, as we can name explicit elements of this restricted subgroup that cannot arise via symmetric monodromy in the parameter space. This stands in direct contrast with reasoning used when studying similar problems, such as in [ACT10, Section 8]. A. Landesman pointed us to the root of this issue: the fibers of our stack quotient used to build the moduli space of symmetric cubic surfaces are not connected.
@@ 1.2. Paper structure
In Section 2, we review the construction of the (marked) moduli space of cubic surfaces, and explicitly realize the moduli space of symmetric cubic surfaces as a GIT quotient. In Section 3 we give an overview of the work of Allcock–Carlson–Toledo, including fundamental facts about framed cubic surfaces and their associated period maps, which yields their main theorem, a uniformization of the moduli of cubic surfaces by complex hyperbolic $4$-space. In Section 4, we analogously define a period map for symmetric cubic surfaces via symmetric framings of cubic surfaces, and uniformize the moduli space of symmetric cubic surfaces by the complex hyperbolic line. This section finishes with a Hodge theoretic restriction of symmetric monodromy in the parameter space to a group of order $96$, which it turns out will properly contain the symmetric monodromy group. In Section 5 we determine that the symmetric monodromy group is the Klein $4$-group $K_{4}$, and we describe how it acts on the $27$ lines. Moreover, we show that the $27$ lines cover over the symmetric locus $\widetilde{\mathcal{Y}}\to\mathcal{Y}$ splits into $12$ connected components, and explicitly determine the $K_{4}$-set structure associated to each connected component. Finally, Section 6 describes how to alternatively witness the symmetric monodromy group $K_{4}$ in $W(E_{6})$ and $\operatorname{PO}(4,1,\mathbb{F}_{3})$ via representation theoretic constructions. Appendix A contains explicit data regarding how the symmetric group $S_{4}$ and the generators of the symmetric monodromy group $K_{4}$ act on the $27$ lines on the Fermat cubic surface.
@@ 1.3. Acknowledgements
We thank Benson Farb, Frank Sottile, and Jesse Wolfson for their interest and all independently asking us what the monodromy group is for the problem of finding lines on symmetric cubic surfaces. We thank Taylor Brysiewicz and Kisun Lee for their help with homotopy continuation and certified tracking software. Thank you to Aaron Landesman, Alberto Landi, and Nick Salter for sharing their insights regarding the relationship between this work and various aspects of big monodromy problems. Lastly, we thank Daniel Allcock for his interest and elucidating answers regarding various group theoretic aspects of his work with James A. Carlson and Domingo Toledo.
@@ 1.4. Notation
$\mathcal{X}$ moduli of anti-canonically embedded cubic surfaces
$\mathcal{Y}$ moduli of anti-canonically embedded symmetric cubic surfaces
$\mathcal{M}$ moduli of cubic surfaces
@@ 2.1. Parameter space of cubic surfaces
There is a linear action of $\operatorname{SL}(4,\mathbb{C})$ on $\mathcal{W}$ given by $g\cdot f:=f\circ g^{-1}$. This induces a left action of $\operatorname{PGL}(4,\mathbb{C})$ on the projectivization $\mathbb{P}(\mathcal{W})$.
@@
The action of $\operatorname{SL}(4,\mathbb{C})$ on each of these loci descends to an action of $\operatorname{PGL}(4,\mathbb{C})$ on their projectivizations. We can take the respective GIT quotients to construct various moduli spaces of cubic surfaces.
@@
When we write a moduli space without a superscript, e.g. $\mathcal{M}$, we implicitly mean the moduli of smooth objects.
The (finite) pointwise stabilizer subgroups in $\operatorname{PGL}(4,\mathbb{C})$ of cubic forms yield automorphisms of the corresponding cubic surface, and thus naturally gives us an orbifold structure on the GIT quotient $\mathcal{M}$. The same will hold true in the symmetric locus, and so we will regard those quotients as orbifolds as well.
@@ 2.2. Marked parameter space of cubic surfaces
It is also classically known that the moduli space of marked cubic surfaces $\widehat{\mathcal{M}}$ is isomorphic to the moduli space of cubic surfaces equipped with six ordered skew lines [Bea09, pg. 19]. We shall freely identify these spaces. Moreover, the cover $\widehat{\mathcal{M}}\to\mathcal{M}$ is the normal closure of the 27 lines cover $\widetilde{\mathcal{M}}\to\mathcal{M}$; we will work on these marked moduli spaces (and their symmetric analogs) to determine our desired Galois groups.
@@ 2.3. Parameter space of symmetric cubic surfaces
We see that the the parameter space of symmetric cubic forms $\mathbb{P}(\mathcal{V})=\mathbb{P}^{2}$ embeds linearly into the parameter space of cubic forms $\mathbb{P}(\mathcal{W})$. Define $\Delta^{S_{4}}$ to be the symmetric discriminant curve, which corresponds to the locus of $S_{4}$-invariant singular cubic forms in the parameter space $\mathcal{Y}$.
@@
The GIT quotient $\mathcal{S}=N_{\operatorname{PGL}(4,\mathbb{C})}(S_{4})\backslash\!\!\backslash\mathcal{Y}^{\mathrm{sm}}$ is the moduli space of smooth symmetric cubic surfaces.
@@
Suppose that two symmetric cubic forms $f_{1},f_{2}\in\mathcal{Y}^{\mathrm{sm}}$ determine isomorphic symmetric cubic surfaces $X=Z(f_{1})$ and $Y=Z(f_{2})$. Since such an isomorphism $\varphi:X\xrightarrow[]{\sim}Y$ of varieties preserves their respective canonical classes $K_{X}$ and $K_{Y}$, the map extends to respect their anticanonical embeddings into $\mathbb{P}^{3}$. Thus such an isomorphism $\varphi$ must be the restriction of a linear automorphism coming from the ambient projective space $\mathbb{P}^{3}$. Moreover, the automorphism groups of $X$ and $Y$ must be preserved under such an isomorphism, and so the $S_{4}$-action on $X$ must be sent to the $S_{4}$-action on $Y$. Thus two symmetric cubic surfaces are projectively equivalent when their symmetric cubic forms differ by an element of the normalizer of $S_{4}<\operatorname{PGL}(4,\mathbb{C})$, which was explicitly calculated in 2.10. ∎
@@ 2.4. Symmetry and stability
The inclusion $\mathcal{V}\to\mathcal{W}$ of parameter spaces descends to a inclusion $\iota:\mathcal{S}\to\mathcal{M}$ of the moduli of symmetric cubic surfaces into the whole moduli space. The pullback $\iota^{*}\widehat{\mathcal{M}}$ is then the moduli space of smooth marked symmetric cubic surfaces $\widehat{\mathcal{S}}$. Then just as before, the pullback construction let’s us conclude that $\iota^{*}\widehat{\mathcal{M}}$ is isomorphic to the moduli space of symmetric cubic surfaces equipped with six ordered skew lines, which we will also denote by $\widehat{\mathcal{S}}$.
@@ 3.4. The period mapping
By 3.3, we have that $\hat{\Gamma}/\hat{\Gamma}^{\prime}\cong W(E_{6})$, and so the Galois cover of $\mathcal{M}$ that $\hat{\Gamma}^{\prime}<\hat{\Gamma}$ corresponds to is the space of cubic surfaces equipped with a line $\widehat{\mathcal{M}}$. Once we have analogously uniformized the moduli space of symmetric cubic surfaces, a special symmetric subgroup of $\mathrm{PO}(4,1,\mathbb{F}_{3})$ will be determined that is the monodromy group of the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$.
@@ 4.3. The global period map
There is an isomorphism of groups $N_{\mathrm{U}(4,1)}(S_{4})\cong\mathrm{U}(1,1)\times(\mathrm{U}(1)\cdot S_{4})$.
@@
Schur’s lemma tells us that the $S_{4}$-centralizer acts by scalars on $W$, and thus is isomorphic to a copy of $\mathrm{U}(1)$ acting on $W$. On the signature $(1,1)$-factor, the $S_{4}$-action is trivial, and thus every element of $\mathrm{U}(1,1)$ arises at an automorphism of the representation $\mathbb{C}\oplus\mathbb{C}$. Thus the normalizer is the product of the normalizers on each factor, proving that $N_{\mathrm{U}(4,1)}(S_{4})\cong\mathrm{U}(1,1)\times(\mathrm{U}(1)\cdot S_{4})$. ∎
@@
There is an isomorphism of groups $N_{\mathrm{U}(1)\times\mathrm{U}(4)}(S_{4})\cong\mathrm{U}(1)\times(\mathrm{U}(1)\cdot S_{4})$, where the second $\mathrm{U}(1)$ acts on the permutation representation $V$ by scalars.
@@
The group $\Gamma=N_{\mathrm{U}(4,1,\mathbb{Z}[\omega])}(S_{4})$ is naturally an arithmetic subgroup of $N_{\mathrm{U}(4,1)}(S_{4})$. Moreover, we have an isomorphism $\Gamma\cong\operatorname{Aut}(\mathrm{diag}(4,-1),\mathbb{Z}[\omega])\times(\langle-\omega\rangle\cdot S_{4})$
@@
Let $\hat{G}=\mathrm{U}(4,1)$, $\hat{K}=\mathrm{U}(4)\times\mathrm{U}(1)$, and $\hat{\Gamma}=\mathrm{U}(4,1,\mathbb{Z}[\omega])$. Set $G=N_{\hat{G}}(S_{4})$, $K=N_{\hat{K}}(S_{4})$, and $\Gamma=N_{\hat{\Gamma}}(S_{4})$. The holomorphic totally geodesic embedding $G/K\to\hat{G}/\hat{K}$ descends to a generically injective finite normalization
@@
Now that we have successfully uniformized the moduli space of symmetric cubic surfaces, we will begin our study of the monodromy group associated to the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$, where $\widetilde{S}$ is the moduli of symmetric cubic surfaces equipped with a line. The following calculation implies Theorem 1.2:
@@
Proof of Theorem 1.2.
Since the inclusion $\mathcal{S}\to\mathcal{M}$ induces an injection on orbifold fundamental groups $P\Gamma\to P\hat{\Gamma}$, it suffices to determine the image of $\Gamma$ in $\operatorname{PO}(4,1,\mathbb{F}_{3})$. This is carried out in 4.12, and is isomorphic to $K_{4}\times S_{4}$. ∎
One may be tempted to conclude that the monodromy group of the cover of parameter spaces $\widetilde{\mathcal{Y}^{\mathrm{sm}}}\to\mathcal{Y}^{\mathrm{sm}}$ is $K_{4}\times S_{4}$. Indeed, [ACT10, Section 8] outlines why the monodromy groups associated to the connected components of moduli of real projective cubic surfaces are the image of their fundamental groups in $\operatorname{PO}(4,1,\mathbb{F}_{3})$. However, all that 4.12 guarantees is that the monodromy group is contained in $K_{4}\times S_{4}$. Remarkably, this fails to pin down our desired Galois group from purely Hodge-theoretic considerations — it will be a proper subgroup of $K_{4}\times S_{4}$! Further analysis using equivariant line geometry on cubic surfaces is required, which we carry out in the next section.